### v3

# TRAIL: Trust-Aware Client Scheduling for Semi-Decentralized Federated Learning

Abstract
Due to the sensitivity of data, Federated Learning (FL) is employed to enable distributed machine learning while safeguarding data privacy and accommodating the requirements of various devices. However, in the context of semi-decentralized FL, clients’ communication and training states are dynamic. This variability arises from local training fluctuations, heterogeneous data distributions, and intermittent client participation. Most existing studies primarily focus on stable client states, neglecting the dynamic challenges inherent in real-world scenarios. To tackle this issue, we propose a TRust-Aware clIent scheduLing mechanism called TRAIL, which assesses client states and contributions, enhancing model training efficiency through selective client participation. We focus on a semi-decentralized FL framework where edge servers and clients train a shared global model using unreliable intra-cluster model aggregation and inter-cluster model consensus. First, we propose an adaptive hidden semi-Markov model to estimate clients’ communication states and contributions. Next, we address a client-server association optimization problem to minimize global training loss. Using convergence analysis, we propose a greedy client scheduling algorithm. Finally, our experiments conducted on real-world datasets demonstrate that TRAIL outperforms state-of-the-art baselines, achieving an improvement of 8.7% in test accuracy and a reduction of 15.3% in training loss.
Introduction
The integration of advanced communication technologies with industrial manufacturing significantly enhances production efficiency and flexibility, accelerating the transition to smart manufacturing (Chen et al. 2024; Lu et al. 2021; Tan et al. 2023). This integration facilitates seamless connectivity between devices and systems through real-time data collection and analysis, which greatly improves the transparency and controllability of production processes (Yu et al. 2023; Wang et al. 2020). Additionally, the incorporation of Artificial Intelligence (AI) further enhances these capabilities. By enabling systems to process and analyze large volumes of data, AI provides solutions for predictive maintenance, intelligent decision-making, and process optimization (Wu et al. 2024).
In modern AI systems, local data on end devices often contains sensitive or private information, rendering traditional edge AI training architectures impractical (Zhang et al. 2024; Wang et al. 2024a). To address security and privacy concerns while minimizing communication costs, a new distributed machine learning framework called Federated Learning (FL) has emerged (Wu et al. 2023; McMahan et al. 2017; Wang et al. 2024b). In FL, each client uploads only model parameters, safeguarding her local data. Typically, this process involves coordination with a single edge server, which can result in high communication overhead and potential single points of failure, particularly in environments with numerous end devices (Zhang et al. 2023a; Lu et al. 2022).
This paper investigates semi-decentralized FL (SD-FL) as a framework to enhance the reliability of model training. As illustrated in Figure 1, we focus on a multi-edge server and multi-client SD-FL framework (Sun et al. 2021). This framework utilizes a two-tier aggregation approach. The first tier is intra-cluster aggregation, where local models are aggregated by their respective servers. The second tier is inter-cluster consensus, which involves exchanging models from multiple servers and collaboratively aggregating them to train a shared global model. By distributing computational and communication loads, this approach enhances both the robustness and scalability of the FL process. While SD-FL mitigates risks associated with single points of failure, existing research often overlooks the dynamic nature of clients, particularly fluctuations in model contributions and communication quality, which can adversely affect training efficiency (Sun et al. 2023).
Research was conducted to address the issue of unreliable clients. In (Sefati and Navimipour 2021), the authors introduced an effective service composition mechanism based on a hidden Markov model (HMM) and ant colony optimization to tackle IoT service composition challenges related to Quality-of-Service parameters. This approach achieved significant improvements in availability, response time, cost, reliability, and energy consumption. In (Ma et al. 2021), the FedClamp algorithm was proposed, which enhanced the performance of the global model in FL environments by utilizing HMM to identify and isolate anomalous nodes. This algorithm was specifically tested on short-term energy forecasting problems. The authors of (Vono et al. 2021) presented the Quantized Langevin Stochastic Dynamics (QLSD) algorithm, which employed Markov Chain Monte Carlo methods to improve dynamic prediction capabilities in FL while addressing challenges related to privacy, communication overhead, and statistical heterogeneity. Additionally, Wang et al. introduced a trust-Age of Information (AoI) aware joint design scheme (TACS) aimed at enhancing control performance and reliability in wireless communication networks within edge-enabled Industrial Internet of Things (IIoT) systems operating in harsh environments. This scheme utilized a learning-based trust model and scheduling strategy. While these studies explored various dynamic aspects, they did not adequately address the interplay between dynamics and client selection strategies.
To bridge this gap, we propose an adaptive hidden semi-Markov model (AHSMM) to predict dynamic changes in training quality and communication quality. AHSMM enhances standard HMM by explicitly modeling state duration distributions, reducing computational complexity, and adapting to dynamic, multi-parameter environments, making it ideal for complex scenarios. To improve SD-FL systems’ control and reliability, we propose a joint mechanism combining dynamic prediction with client selection. Extensive experiments and analyses validate the effectiveness and robustness of our approach under varying client dynamics. The main contributions of this paper are as follows:
  • We propose a unified optimization mechanism named TRAIL for SD-FL that integrates performance prediction and client scheduling, enhancing model robustness, accelerating convergence speed, and improving overall performance.
  • We introduce an AHSMM to predict client performance and channel variations to obtain trust levels. This model effectively accounts for both dynamic and static aspects of clients, enabling efficient state predictions for each one.
  • Through convergence analysis, we assess the anticipated effects of client-server relationships on convergence. This analysis allows us to reformulate the initial optimization challenge as an integer nonlinear programming problem, for which we devise a greedy algorithm to optimize client scheduling efficiently.
  • Extensive experiments conducted on four real-world datasets (MNIST, EMNIST, CIFAR10, SVHN) demonstrate that our proposed mechanism outperforms state-of-the-art baselines, achieving an 8.7% increase in test accuracy and a 15.3% reduction in training loss.
Related Work
In FL, model training is distributed across multiple clients to protect data privacy and minimize the need for centralized data aggregation. Traditional FL assumed reliable and frequent communication between clients and the server. However, this assumption often failed in real-world applications, particularly in environments with heterogeneous devices and unstable communication. To address these challenges, researchers introduced SD-FL. This framework combined the benefits of centralized and distributed architectures by enabling direct communication among some clients, thereby reducing the server’s workload and communication costs. SD-FL was better equipped to adapt to dynamic network environments and heterogeneous data distributions, enhancing the system’s robustness and efficiency.
Research efforts primarily focused on the following two areas. (i) Client Selection: Mechanisms were developed to select clients for participation in training, ensuring that chosen clients met performance criteria, thereby enhancing the overall effectiveness of the model (Lin et al. 2021; Wang et al. 2024c; Yemini et al. 2022; Sun et al. 2023). (ii) Trust Management: Trust mechanisms were introduced to assess and predict the reliability of clients, ensuring that only reliable clients participated in training. These mechanisms contributed to improved model robustness and performance (Martínez Beltrán et al. 2023; Parasnis et al. 2023; Xu et al. 2024; Valdeira et al. 2023).
While some studies focus separately on client selection and trust management, existing methods often fail to effectively integrate these aspects, which negatively impacts the efficiency and performance of SD-FL systems. Our work proposes an integrated mechanism that addresses client selection and trust management, which is essential for enhancing the robustness and performance of SD-FL, especially in diverse and potentially unreliable environments.
System Model
Figure 1: The SD-FL system framework.
Here, we explore the SD-FL framework, as illustrated in Figure 1. We first present SD-FL’s basic workflow, then establish an adaptive semi-Markov model to estimate each client’s model quality and communication quality.
Basics of SD-FL
We examine the SD-FL training process across $T$ rounds, involving $S$ edge servers represented by $\mathcal{S}=\{1,2,\cdots,S\}$, and $U$ client devices represented by $\mathcal{U}=\{1,2,\cdots,U\}$. Each round consists of the following steps:
  • Clients perform $T_{1}$ rounds of local training using their datasets. Then, they upload the trained local models to the edge server for intra-cluster aggregation.
  • After aggregating the models at the edge server, the merged model is broadcasted to the corresponding clients for model updating.
  • After $T_{2}$ rounds of intra-cluster aggregation, each edge server sends its latest model to neighboring servers to achieve inter-cluster consensus.
  • After aggregating to obtain inter-cluster models, these models are sent back to their respective clients for the next round of training.
AHSMM Model
The Adaptive Hidden Semi-Markov Model (AHSMM) extends the traditional Hidden Semi-Markov Model (HSMM) (McDonald, Gales, and Agarwal 2023) into adaptive training using multi-parameter information (i.e., client training accuracy, packet loss), thereby enhancing both modeling and analytical capabilities. The AHSMM model can be described by the parameters $\Sigma=(\pi,A,B,E)$, where: $\pi$ represents the initial state probabilities, $A$ denotes the macro state transition probabilities, $B$ corresponds to the observation probabilities after adaptive training, $E$ represents the state dwell time after adaptive training, encompassing both the existing and remaining dwell times. In addition, similar to HSMM, AHSMM addresses three core problems: evaluation, recognition, and training. To this end, AHSMM defines new forward-backward variables and proposes improved algorithms for forward-backward processes, the Viterbi algorithm (Zhang et al. 2023b), and the Baum-Welch algorithm (Zhang et al. 2023c).
The computational complexity of the Hidden Semi-Markov Model (HSMM) is relatively high. To address this complexity, the Adaptive Hidden Semi-Markov Model (AHSMM) introduces a new forward variable, denoted as $\alpha_{t}(i,e)$. This variable represents the probability of generating the observations $z_{1},z_{2},\ldots,z_{t}$, given that the quality state $i$ has a specific dwell time of $e_{t}(i,e)=e$. In this context, $\varepsilon_{t}$ signifies the current dwell time of the quality state $q_{t}$. When $\left(q_{t},\varepsilon_{t}\right)$ is assigned the value $(i,e)$, it indicates that the device has remained in its current quality state $i$ up to time $t$. During this period, the state $i$ has accumulated a dwell time of $e$ and is prepared to transition to a different quality state at time $t+$ 1. Therefore, for $11\leq t\leq T-$ 1 and $e\in[1,E]$, we can define the forward variable as follows:
$\alpha_{t}(i,e)=p\left(z_{1}^{t},\left(q_{[t-e+1,t]}=i,\varepsilon_{t}=e\right)\mid\Sigma\right).$ (1)
The forward recursion is obtained as:
$\alpha_{t}(i,e)=\begin{cases}\sum_{j\neq i}^{N}a_{ji}b_{i}(z_{t})\left(\sum_{\tau=1}^{E}\alpha_{t-1}(j,\tau)p_{j}(\tau)\right),&\\ \hskip 150.00023pt\text{if }e=1,\\ \alpha_{t-1}(i,e-1)\prod_{s=1}^{e}b_{i}(z_{t-s+1}),&\\ \hskip 150.00023pt\text{if }e>1,\\ \pi_{i}b_{i}(z_{1})p_{i}(e),&\\ \hskip 150.00023pt\text{if }t=1,\\ 0,&\\ \hskip 150.00023pt\text{if }\tau<1.\end{cases}$ (2)
In the context of AHSMM, let $E$ represent the maximum state dwell time among all quality states. Given the model $\Sigma$, the probability of observing the sequence $Z$ is expressed as:
$p\left(z_{1}^{T}\mid\Sigma\right)=\sum_{(i,e)}\alpha_{T}(i,e).$ (3)
The variable $\alpha_{T}(i,e)$ is defined as the joint probability of observing the sequence $z_{1},z_{2},\ldots,z_{T}$ while the system is in quality state $i$ over the dwell time interval from $T-e+1$ to $T$. Mathematically, it can be expressed as:
$\alpha_{T}(i,e)$ $=p\left(q_{[T-e+1,T]}=i\mid z_{1},z_{2},\ldots,z_{T},\Sigma\right)$ (4)
$\cdot\,p\left(z_{1},z_{2},\ldots,z_{T}\mid\Sigma\right).$
For $11\leq t\leq T-$ 1, $e\in[1,E]$, and $i,j\in S$, the backward variable can be defined as:
$\beta_{t}(i,e)=p\left(z_{(t+1):T}\mid q_{[t-e+1,t]}=i,\Sigma\right).$ (5)
This formulation improves the efficiency of computing the forward and backward variables in the AHSMM, leading to reduced computational complexity compared to the traditional HSMM. In the backward variable $\beta_{t}(i,e)$, the quality state $i$ has been active for $e$ time steps. By summing over all quality states and potential dwell times, the backward recursion can be expressed as follows:
$\beta_{t}(i,e)=\sum_{j}\left(\mathbb{I}(j\neq i)\beta_{t+1}(j,1)a_{ij}p_{i}(e)b_{j}\left(z_{t+1}\right)\right)$ (6)
$+\beta_{t+1}(j,e+1)\prod_{s=1}^{e}b_{i}\left(z_{t+s}\right).$
We estimate the quality state $q_{t}$ and update the parameters of the model $\Sigma$. Using the previously defined forward and backward variables, we derive $q_{t}$ and adjust the model parameters. Given the model $\Sigma$ and the observation sequence $Z_{1:T}$, let $\xi_{t}^{e}(i,j)$ represent the joint probability of the observation sequence $Z_{1:T}$ and the transition from quality state $i$ to quality state $j$ (where $i\neq j$ ) at time $t$. The specific formula is as follows:
$\xi_{t}^{e}(i,j)=p\left(z_{1:T}\mid q_{t-1}=i,q_{t}=j,\Sigma\right)$ (7)
$\cdot p\left(q_{t-1}=i,q_{t}=j\mid\Sigma\right).$
To determine quality states from a sequence of observations, it is essential to have both a predefined model and the sequence of observations. The following equation describes the recursive estimation of quality states based on this model:
$\gamma_{t}^{e}(i)=\sum_{j=1}^{E}\alpha_{t}(j,e)\cdot p\left(q_{t}=i\mid\Sigma\right)$ (8)
$\cdot p\left(z_{1:T}\mid q_{t}=i,\Sigma\right).$
This recursive formulation allows for the estimation of the quality state at time $t$ based on the observation sequence $Z_{1:T}$ and the given model parameters $\Sigma$.
AHSMM Prediction and Client Scheduling
AHSMM Parameter Estimation
Monitoring a device with multiple parameters can significantly improve quality prediction. Given the inherent differences among parameters, effective data fusion is essential for integrating their information. Consequently, estimating the parameters of the AHSMM becomes necessary. This estimation process utilizes Maximum Likelihood Linear Regression (MLLR) transformations to address the variations across parameters. Simultaneously, a canonical model is trained based on a set of MLLR transformations. Linear transformations are then applied to the mean vectors of the state output and dwell time distributions in the standard model, allowing for the derivation of mean vectors for these distributions. The formulas are given by:
$\begin{cases}b_{i}\left(z^{(s)}\right)=N\left(Z;\mu_{i}^{(s)}\Sigma_{i}\right),\\ p_{i}(e)=N\left(e;\mu_{e}^{(s)},\sigma_{i}^{2}\right),\quad\\ \mu_{e}^{(s)}=\delta^{(s)}m_{i}+\psi^{(s)}.\end{cases}.$ (9)
Here, $b_{i}\left(\boldsymbol{z}^{(s)}\right)$ represents the probability density function for the state $i$ based on the observed data $z^{(s)}$ from parameter $s$, modeled as a multivariate normal distribution with mean $\mu_{i}^{(s)}$ and covariance $\Sigma_{i}$. The term $p_{i}(e)$ signifies the probability density function for the dwell time $e$ in state $i$, also following a normal distribution characterized by mean $\mu_{e}^{(s)}$ and variance $\sigma_{i}^{2}$. The mean dwell time $\mu_{e}^{(s)}$ is calculated using the formula $\mu_{e}^{(s)}=\delta^{(s)}m_{i}+\psi^{(s)}$, where $\delta^{(s)}$ is a scaling factor, $m_{i}$ represents a parameter related to state $i$, and $\psi^{(s)}$ is a parameter-specific offset.
Let $S$ denote the number of parameters, and let $\boldsymbol{Z}=(\boldsymbol{Z}^{(1)},\cdots,\boldsymbol{Z}^{(S)})$ represent the monitoring data, where $\boldsymbol{Z}^{(s)}=(z_{1s},\cdots,z_{Ts})$ represents the monitoring data of parameter $s$ with length $T_{s}$. Here, the parameters are estimated by jointly considering the contributions of all parameters and their respective transformations. The term $\gamma_{t}^{e}(i)$ represents the probability of being in state $i$ with dwell time $d$ at time $t$, $\eta^{(s)}$ and $\xi^{(s)}$ are the transformation matrices and vectors for parameter $s$, and $\Sigma_{i}$ is the covariance matrix for state $i$. This joint estimation process ensures that the model parameters are optimally adjusted for the diverse parameter data. The MAP estimation of state $q_{t}$ using the AHSMM is calculated in the following way:
$\hat{q}_{t}=\arg\max_{\Sigma}\prod_{s=1}^{S}\left(p\left(Z^{s}\mid q_{t},\Sigma\right)\right..$ (10)
In client quality diagnostics and forecasting, parameters like computation, communication, and data quality influence decision-making differently. The AHSMM effectively integrates these diverse parameters by capturing temporal dependencies and assigning appropriate weights. This enables accurate client quality assessment, improving forecasting and scheduling in dynamic environments.
Prediction Process Based on AHSMM
Assuming $F(0)=0$ and the failure probability density function $f(t)=F^{\prime}(t)$, the HR function is defined as:
$\Sigma(t)=\frac{f(t)}{1-F(t)}=\frac{\mathrm{d}k(t)}{N-k(t)\mathrm{d}t}.$ (11)
A device transitions through multiple quality states before ultimately reaching a failure state. Let $E(i)$ represent the residence time in quality state $i$. We can express $E(i)$ as follows:
$E(i)=m(i)+\rho\sigma^{2}(i),$
where $m(i)$ denotes the mean dwell time in state $i$ and $\sigma^{2}(i)$ is the variance of the dwell time in that state, and the term $\rho$ serves as a proportionality constant, which adjusts the influence of the variance on the overall residence time.
This formulation captures the idea that the total time spent in a quality state is influenced not only by the average time spent there but also by the variability of that time. A higher variance indicates greater uncertainty in the duration spent in state $i$, which can lead to longer overall residence times. By incorporating both mean and variance, we obtain a more comprehensive view of the dynamics in quality states. The proportionality constant $\rho$ is defined as:
$\rho=\frac{T-\sum_{i=1}^{N}m(i)}{\sum_{i=1}^{N}\sigma^{2}(i)},$ (12)
where $T$ is the total lifespan, $m(i)$ is the mean dwell time in state $i$, and $\sigma^{2}(i)$ is the variance.
The reliability function $R(t+e\Delta t)$ represents the probability that the client remains in the current quality state $i$ at time $t+e\Delta t$. Thus, we can get
$\gamma_{t}^{e}(i)=R(t+e\Delta t).$ (13)
and
$\frac{\bar{\Sigma}(t+e\Delta t)}{\gamma_{t}^{e}(i)}=\frac{\xi_{t}^{e}(i,j)}{\Delta t}.$ (14)
Based on the above equations, $\bar{E}(i,e)$ can be expressed as:
$\bar{E}(i,e)=E(i)-\frac{E(i)\xi_{t}^{e}(i,j)}{\gamma_{t}^{e}(i)}.$ (15)
According to equations (13,14,15), once the client reaches state $i$ and has a dwell time, the trust level (TL) is determined as:
$TL_{(i,e)}=\bar{E}(i,e)+\left(\sum_{j=1}^{N}E(j)-\sum_{j=1}^{i}E(j)\right).$ (16)
Client Selection
Considering the configuration in SD-FL, the model aggregation within the cluster at server $s$ can be described as follows:
$\boldsymbol{w}_{s,t}=\sum_{i\in U_{s}}\frac{n_{i}}{\sum n_{i}}\boldsymbol{w}_{i,t},$ (17)
where $U_{s}$ represents the set of clients assigned to server $s$, $n_{i}$ is the size of the local dataset for client $i$, and $\boldsymbol{w}_{i,t}$ denotes the local model of client $i$ at training round $t$. This weighted aggregation ensures that clients with larger datasets contribute proportionally more to the cluster model, improving the robustness of the overall learning process.
Furthermore, the model consensus between clusters at server $s$ during training round $t$ is characterized as follows:
$\boldsymbol{g}_{s,t}=\sum_{s=1}^{S}\frac{N_{s}}{\sum N_{s}}\boldsymbol{w}_{s,t},$ (18)
where $N_{s}$ corresponds to the total size of the datasets managed by server $s$, and $S$ is the number of servers. This global aggregation step aligns the models from different clusters, ensuring consistency and convergence across the distributed system.
In the subsequent training round, each client utilizes the updated global model $\boldsymbol{g}_{s,t}$ as the initialization for their local model updates. Clients then train their local models using their respective datasets through the gradient descent mechanism, defined as follows:
$\boldsymbol{w}_{i,t+1}=\boldsymbol{w}{i,t}-\eta\nabla F_{i}(\boldsymbol{w}_{i,t}),$ (19)
where $\eta$ is the learning rate, and $\nabla F_{i}(\boldsymbol{w}_{i,t})$ represents the gradient of the local loss function $F_{i}$ with respect to the current model $\boldsymbol{w}_{i,t}$.
This decentralized training mechanism enables clients to collaboratively train a global model while keeping their data local, thereby addressing privacy concerns and minimizing the communication overhead associated with transmitting raw data. The combination of local training, cluster-level aggregation, and global consensus helps achieve a balance between computational efficiency, communication cost, and model accuracy in distributed learning systems.
Problem Formulation
Client scheduling determines the optimal client-server association matrix $\boldsymbol{d}$ to minimize the global model loss. Each element $d_{ij}$ is binary (1 if client $i$ is assigned to server $j$, 0 otherwise). The configuration of $\boldsymbol{d}$ directly impacts the global loss by influencing data locality, communication overhead, and computational load balancing. Adaptive scheduling dynamically adjusts $\boldsymbol{d}$ to further enhance system performance and ensure efficient training. The global loss function is defined as:
$F(\boldsymbol{g})=\frac{1}{N}\sum_{s=1}^{S}\sum_{i\in\mathcal{U}_{s}}F_{i}\left(\boldsymbol{g}_{s}\right).$ (20)
Therefore, the optimization of the client-server association matrix can be achieved by solving the problem of minimizing the global training loss:
$\min_{\boldsymbol{d}}$ $\quad F(\boldsymbol{g})=\frac{1}{N}\sum_{s=1}^{S}\sum_{i\in\mathcal{U}_{s}}F_{i}\left(\boldsymbol{g}_{s}\right)$ (21)
subject to $\quad\sum_{\forall s\in\mathcal{S}}d_{i,s}\leq 1,$
$\quad d_{i,s}\in\{0,1\},$
$\quad TL_{i,s}\geq\Theta.$
Convergence Analysis
We rely on current trust levels of client $i$ to tackle these challenges, reduce data loss, and guarantee consistent model updates. $\Theta$ denotes the threshold for the trust level of clients involved in training. Reformulating the optimization problem to incorporate parameters that reflect communication link stability will enhance the modeling of the SD-FL system’s conditions.
We also plan to enhance SD-FL system robustness through redundancy strategies, like multiple communication paths or backup servers, to mitigate risks from unreliable links. Dynamically adjusting client-server associations based on real-time assessments will help maintain optimal performance despite trust fluctuations. This approach maximizes resource utilization and minimizes training time, leading to more robust convergence and broader adoption in real-world applications.
Theorem 1
By setting the learning rate $\lambda=\frac{1}{L}$, the upper bound of the expected difference $\mathbb{E}\left(F\left(\boldsymbol{g}_{t+1}\right)-F\left(\boldsymbol{g}^{*}\right)\right)$ can be established as follows:
[…]
where $D=1-\frac{\mu}{L}+\frac{4\omega_{2}\mu B}{L}$ , $B=\sum_{m\in\mathcal{S}}\frac{1}{N_{m}^{(S)}}\Psi$ , and $\Psi=\left(\sum_{i\in\mathcal{U}}n_{i}d_{i,m}\left(D_{m}-1+\mathbb{I}\left(TL_{i,s}<\Theta\right)\right)\right)$.
In the definition of $B$, the following equations hold:
$N_{m}^{(S)}=\sum_{i\in\mathcal{U}}n_{i}d_{i,m},$ (23)
and
$D_{m}=\frac{1}{1+|S|}.$ (24)
Input: Set of clients $\mathcal{U}$, set of servers $\mathcal{S}$, maximum clients per server $U_{\max}$, the allocation status $x_{i}$;
Output: Assignment matrix $\boldsymbol{d}$;
1 Initialize $\boldsymbol{d}_{i,s}=0,\forall i\in\mathcal{U},\forall s\in\mathcal{S}$;
2 Initialize client count $u_{s}=0,\forall s\in\mathcal{S}$;
3 Sort $(i,s)\in(\mathcal{U},\mathcal{S})$ in increasing order based on $TL_{i,s}$ for each server $s$;
4 for (i,s) in $(\mathcal{U},\mathcal{S})$ do
5       if $u_{s}<U_{\max}$ and $x_{i}=0$ and $TL_{i,s}\geq\Theta$  then
6             Assign client $i$ to server $s$: $d_{i,s}=1$;
7             Increment client count for server $s$: $u_{s}=u_{s}+1$;
8             $x_{i}=1$;
9       end if
10
11 end for
Algorithm 1 Greedy Algorithm for Solving the Optimization Problem
From Theorem 1, the global training loss minimization initially outlined in the problem can be reinterpreted to focus primarily on minimizing the parameter $B$. This revised formulation, therefore, positions $B$ as the central target for reduction, aiming to directly influence and improve the overall system performance by addressing the underlying factors contributing to $B$ ’s value, i.e.,
$\min_{d}$ $\sum_{m\in\mathcal{S}}\frac{1}{N_{m}^{(S)}}\left(\sum_{i\in\mathcal{U}}n_{i}d_{i,m}\left(D_{m}-1+\mathbb{I}\left(TL_{i}<\Theta\right)\right)\right).$ (25)
To address this nonlinear integer programming problem, we propose a greedy algorithm outlined in Algorithm 1, with a time complexity of $O(nm\log nm)$.
Experiments Evaluation
This section evaluates our proposed mechanism using real-world datasets to demonstrate its effectiveness and practicality. We begin by introducing the experiment setting. Then, we present our experimental comparisons’ results, highlighting our mechanism’s performance relative to the baselines.
Experments Setting
We provide a detailed explanation of the fundamental experimental setup, including the basic setup, datasets, training configurations, baselines, and evaluation metrics.
Basic Setup: We design a SD-FL system comprising five edge servers and fifty clients, with each client assigned 1,000 local training samples. To emulate real-world challenges, 10%, 30%, and 50% of the clients gradually experience degradation in both training quality (e.g., training accuracy) and communication quality (e.g., packet loss) as training progresses.
Datasets: Real-world Datasets. Four standard real-world datasets, e.g. MNIST (Rana, Kabir, and Sobur 2023), EMNIST (Majeed et al. 2024), SVHN (Pradhan et al. 2024),and CIFAR-10 (Aslam and Nassif 2023) are utilized to make performance evaluation.
Training Configurations: Training Parameters. We adopt a CNN architecture for its effectiveness in image processing tasks. The batch size is 32, balancing computational efficiency and model performance. Each client performs 100 local training rounds ($T_{1}=100$) before aggregation, with 100 inter-cluster aggregations ($T_{2}=100$) to synchronize updates across edge servers. The learning rate ($\eta$) is set to 0.01, ensuring stable and efficient optimization, and SGD with a momentum of 0.05. The model uses ReLU as the activation function for non-linearity and cross-entropy loss for classification tasks.
Baselines: In order to validate the effectiveness of our proposed mechanism, we compared our mechanism with the following three mechanisms.
  • GUROBI: In (Muley 2021), the authors utilize the GURUBI optimizer for the client’s optimal allocation problem. The prediction part of the front end does not utilize the prediction mechanism of AHSMM.
  • TRUST. In (Wang et al. 2024c), the authors introduce a trust-age of information (AoI)-aware co-design scheme (TACS), employing a learning-based trust model and trust-AoI-aware scheduling to optimize data selection for plant control dynamically.
  • RANDOM. Here, we continue to use the AHSMM for the prediction component, while employing random allocation for client assignments.
Evaluation metrics: We use two metrics to evaluate our mechanisms: test accuracy and training loss. The results are obtained from the average of multiple experiments.
  • Test accuracy. Test accuracy measures a model’s performance on unseen data, reflecting its generalization ability and effectiveness in SD-FL, critical for real-world usability and reliability.
  • Training loss. Training loss quantifies the discrepancy between the predicted outputs of a model and the actual data, guiding the optimization process to improve model accuracy and performance.
Figure 2: The test accuracy and training loss in scenarios with 10% low-quality clients: (a) MNIST, (b) EMNIST, (c) CIFAR10, and (d) SVHN.
Figure 3: The test accuracy in scenarios with 10%,30%,50% low-quality clients: (a) MNIST, (b) EMNIST, (c) CIFAR10, and (d) SVHN.
Experiments Results
In Figure 2, we analyze the variations in test accuracy and training loss over multiple training rounds for four distinct mechanisms, specifically under conditions where only 10% of clients are classified as low quality. Our proposed mechanism stands out by achieving the highest performance across four real datasets. This result is due to the effective integration of our AHSMM model with a greedy algorithm, which works together to predict fluctuations in client learning and communication quality reasonably. By optimizing the participation of low-quality clients, our mechanism significantly enhances overall training outcomes. In contrast, the TRUST mechanism, despite employing a predictive approach, lacks an effective client distribution strategy. This deficiency leads to suboptimal performance, as it fails to adaptively manage client participation based on their quality. Similarly, the RANDOM mechanism incorporates the AHSMM to forecast client behavior, but it does not allocate clients efficiently, leading to less effective training sessions. Although the GUROBI mechanism can determine the optimal client allocation scheme, it does not incorporate client quality predictions, which hampers its ability to promptly exclude low-quality clients from participating in training, ultimately affecting training efficiency. Overall, our mechanism demonstrates a superior ability to navigate the complexities of client quality by leveraging predictive modeling and strategic allocation, ensuring robust training performance in SD-FL environments.
In Figure 3, we analyze the comparative performance of four mechanisms across scenarios characterized by varying proportions of low-quality clients. As the percentage of low-quality clients increases, all mechanisms demonstrate a decline in both training and testing accuracy, albeit to different extents. Notably, our proposed mechanism consistently delivers the best results across four real datasets, showcasing its robustness in challenging environments. The superior performance of our mechanism can be attributed to the integration of the AHSMM and a greedy-based client allocation algorithm. This combination effectively predicts fluctuations in client learning and communication quality, enabling efficient client allocation that significantly enhances model training quality, even under adverse conditions. By optimizing the participation of higher-quality clients, we mitigate the negative impact of low-quality clients on overall performance. In contrast, while the TRUST mechanism is capable of identifying unreliable clients, it lacks an efficient client distribution strategy. This limitation results in poorer training outcomes compared to our mechanism, as it fails to adaptively manage client participation based on their quality. Similarly, the RANDOM mechanism employs the AHSMM to accurately predict changes in client training quality but relies on a random allocation strategy during the client distribution phase. Consequently, this randomness undermines the final training effectiveness, leading to suboptimal results. The GUROBI mechanism, despite its potential for determining optimal client distributions, performs the worst in our experiments. Its inability to accurately predict changes in client learning quality restricts its effectiveness, as it cannot exclude low-quality clients in a timely manner. Experiments reveal a notable 8.7% increase in test accuracy and a 15.3% reduction in training loss compared to existing baselines, demonstrating the superiority of our mechanism in SD-FL settings. These results underscore the importance of both predictive modeling and strategic client allocation in achieving high-quality training outcomes.
Conclusion
This paper proposes TRAIL, a novel mechanism designed to address the dynamic challenges in SD-FL. TRAIL integrates an AHSMM to accurately predict client states and contributions and a greedy algorithm to optimize client-server associations, effectively minimizing global training loss. Through convergence analysis, the impact of client-server relationships on model convergence is theoretically assessed. Extensive experiments conducted on four real-world datasets demonstrate that TRAIL improves test accuracy and training loss, significantly outperforming state-of-the-art baselines. This work highlights the potential of combining predictive modeling and strategic client allocation to enhance efficiency, robustness, and performance in distributed learning systems.
Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grants 62372343, 62072411, and 62402352, in part by the Zhejiang Provincial Natural Science Foundation of China under Grant LR21F020001, in part by the Key Research and Development Program of Hubei Province under Grant 2023BEB024, and in part by the Open Fund of Key Laboratory of Social Computing and Cognitive Intelligence (Dalian University of Technology), Ministry of Education under Grant SCCI2024TB02.
Appendix
Proof of Theorem 1
Assumption 1: Strong Convexity and Lipschitz Continuity Assume the global loss function $\mathrm{F}(g)$ is $\mu$-strongly convex and its gradient is $L$-Lipschitz continuous. This implies that for any $g$ and $g^{\prime}$:
  • Strong convexity:
    $F\left(g^{\prime}\right)\geq F(g)+\left\langle\nabla F(g),g^{\prime}-g\right\rangle+\frac{\mu}{2}\left\|g^{\prime}-g\right\|^{2}.$
  • Gradient Lipschitz continuity:
    $\left\|\nabla F(g)-\nabla F\left(g^{\prime}\right)\right\|\leq L\left\|g-g^{\prime}\right\|.$
Assumption 2: Assume that during training, randomness is introduced (e.g., due to random client participation), and these random factors are independent and identically distributed (i.i.d.).
Definitions: Let $g_{t}$ denote the global model at iteration $t$ , and $g^{*}$ denote the global optimal model. Let $omega_{1}$ and $\omega_{2}$ be constants related to the system, and $B$ is a cumulative error term. In FL, the global model update can be expressed as:
$g_{t+1}=g_{t}-\lambda\nabla F\left(g_{t}\right)+\eta_{t},$
where $\eta_{t}$ represents the noise or error term due to factors like client sampling and unreliable communication. Since $\lambda=\frac{1}{L}$, we have:
$g_{t+1}=g_{t}-\frac{1}{L}\nabla F\left(g_{t}\right)+\eta_{t}.$
Consider the difference:
$F\left(g_{t+1}\right)-F\left(g^{*}\right).$
Using the $L$-Lipschitz continuity of the gradient, we have:
$F\left(g_{t+1}\right)\leq F\left(g_{t}\right)+\left\langle\nabla F\left(g_{t}\right),g_{t+1}-g_{t}\right\rangle+\frac{L}{2}\left\|g_{t+1}-g_{t}\right\|^{2}$
Substituting $9_{\{}t+1\}-g_{t}$, we get:
$g_{t+1}-g_{t}=-\frac{1}{L}\nabla F\left(g_{t}\right)+\eta_{t},$
so
$\left\|g_{t+1}-g_{t}\right\|\leq\frac{1}{L}\left\|\nabla F\left(g_{t}\right)\right\|+\left\|\eta_{t}\right\|.$
Substitute back into the inequality:
$F\left(g_{t+1}\right)\leq F\left(g_{t}\right)-\frac{1}{L}\left\|\nabla F\left(g_{t}\right)\right\|^{2}+\left\langle\nabla F\left(g_{t}\right),\eta_{t}\right\rangle$
$+\frac{L}{2}\left(\frac{1}{L}\left\|\nabla F\left(g_{t}\right)\right\|+\left\|\eta_{t}\right\|\right)^{2}.$
Take the expectation of both sides. Assuming that $\eta_{t}$ has zero mean (the noise is mean zero), we have:
$E\left[F\left(g_{t+1}\right)\right]\leq E\left[F\left(g_{t}\right)\right]-\frac{1}{L}E\left[\left\|\nabla F\left(g_{t}\right)\right\|^{2}\right]+$
$\frac{L}{2}E\left[\left(\frac{1}{L}\left\|\nabla F\left(g_{t}\right)\right\|+\left\|\eta_{t}\right\|\right)^{2}\right].$
Expanding the squared term:
$\left(\frac{1}{L}\left\|\nabla F\left(g_{t}\right)\right\|+\left\|\eta_{t}\right\|\right)^{2}=\left(\frac{1}{L}\left\|\nabla F\left(g_{t}\right)\right\|\right)^{2}+$
$2\frac{1}{L}\left\|\nabla F\left(g_{t}\right)\right\|\left\|\eta_{t}\right\|+\left\|\eta_{t}\right\|^{2}.$
Therefore, we can get
$E\left[F\left(g_{t+1}\right)\right]\leq E\left[F\left(g_{t}\right)\right]-$
$\left(\frac{1}{L}-\frac{1}{2L}\right)E\left[\left\|\nabla F\left(g_{t}\right)\right\|^{2}\right]+$
$LE\left[\left\|\nabla F\left(g_{t}\right)\right\|\left\|\eta_{t}\right\|\right]+\frac{L}{2}E\left[\left\|\eta_{t}\right\|^{2}\right].$
Note that $\frac{1}{L}-\frac{1}{2L}=\frac{1}{2L}$. Thus,
$E\left[F\left(g_{t+1}\right)\right]\leq E\left[F\left(g_{t}\right)\right]-\frac{1}{2L}E\left[\left\|\nabla F\left(g_{t}\right)\right\|^{2}\right]+$
$LE\left[\left\|\nabla F\left(g_{t}\right)\right\|\left\|\eta_{t}\right\|\right]+\frac{L}{2}E\left[\left\|\eta_{t}\right\|^{2}\right].$
Since $F(g)$ is $\mu$-strongly convex, we have:
$\left\|\nabla F\left(g_{t}\right)\right\|^{2}\geq 2\mu\left(F\left(g_{t}\right)-F\left(g^{*}\right)\right)$
$E\left[F\left(g_{t+1}\right)\right]\leq E\left[F\left(g_{t}\right)\right]-\frac{\mu}{L}E\left[F\left(g_{t}\right)-F\left(g^{*}\right)\right]+$
$LE\left[\left\|\nabla F\left(g_{t}\right)\right\|\left\|\eta_{t}\right\|\right]+\frac{L}{2}E\left[\left\|\eta_{t}\right\|^{2}\right.$
Introduce constants $\omega_{1}$ and $\omega_{2}$ satisfying:
$E\left[\left\|\nabla F\left(g_{t}\right)\right\|\left\|\eta_{t}\right\|\right]\leq\omega_{1}E\left[F\left(g_{t}\right)-F\left(g^{*}\right)\right]+\omega_{2}B,$
and
$E\left[\left\|\eta_{t}\right\|^{2}\right]\leq B$
where $B$ is the cumulative error term defined earlier. Substitute these into the inequality:
$E\left[F\left(g_{t+1}\right)\right]\leq E\left[F\left(g_{t}\right)\right]-$
$\left(\frac{\mu}{L}-L\omega_{1}\right)E\left[F\left(g_{t}\right)-F\left(g^{*}\right)\right]+\left(L\omega_{2}+\frac{L}{2}\right)B.$
Thus,
$E\left[F\left(g_{t+1}\right)-F\left(g^{*}\right)\right]\leq$
$\left(1-\frac{\mu}{L}+L\omega_{1}\right)E\left[F\left(g_{t}\right)-F\left(g^{*}\right)\right]+\left(L\omega_{2}+\frac{L}{2}\right)B.$
Note that $D=1-\frac{\mu}{L}+\frac{4\omega_{2}\mu B}{L}$. We can adjust constants $\omega_{1}$ and $\omega_{2}$ (since they are related to $B$) to satisfy:
$L\omega_{1}=\frac{4\omega_{2}\mu B}{L}.$
Therefore, the inequality becomes:
$E\left[F\left(g_{t+1}\right)-F\left(g^{*}\right)\right]\leq DE\left[F\left(g_{t}\right)-F\left(g^{*}\right)\right]+\frac{2\omega_{1}B}{L}.$
By iteratively applying the inequality from $t=0$ to $t$, we get:
$E\left[F\left(g_{t+1}\right)-F\left(g^{*}\right)\right]\leq D^{t+1}E\left[F\left(g_{0}\right)-F\left(g^{*}\right)\right]+$
$\frac{2\omega_{1}B}{L}\sum_{k=0}^{t}D^{k}.$
Using the geometric series formula:
$\sum_{k=0}^{t}D^{k}=\frac{1-D^{t+1}}{1-D}$
we have:
$E\left[F\left(g_{t+1}\right)-F\left(g^{*}\right)\right]\leq D^{t+1}E\left[F\left(g_{0}\right)-F\left(g^{*}\right)\right]+$
$\frac{2\omega_{1}B}{L}\cdot\frac{1-D^{t+1}}{1-D}.$
To align with the statement of Theorem 1, we replace $t+1$ with $t$, so the final result is:
$E\left[F\left(g_{t}\right)-F\left(g^{*}\right)\right]\leq D^{t}E\left[F\left(g_{0}\right)-F\left(g^{*}\right)\right]+\frac{2\omega_{1}B}{L}\cdot\frac{1-D^{t}}{1-D}.$
We have thus proven that for the learning rate $\lambda=\frac{1}{L}$, the expected difference in function values satisfies:
$E\left[F\left(g_{t}\right)-F\left(g^{*}\right)\right]\leq D^{t}E\left[F\left(g_{0}\right)-F\left(g^{*}\right)\right]+\frac{2\omega_{1}B}{L}\cdot\frac{1-D^{t}}{1-D},$
where
$D=1-\frac{\mu}{L}+\frac{4\omega_{2}\mu B}{L}.$
In the above proof, $B$ is the cumulative error term, which is related to factors like client selection and communication quality. Specifically, $B$ is defined as:
$B=\sum_{m\in\mathcal{S}}\frac{1}{N_{m}^{(S)}}\Psi$
where
$\Psi=\left[\sum_{i\in\mathcal{U}}n_{i}d_{i,m}\left(D_{m}-1+I\left(TL_{i,s}<\Theta\right)\right)\right],$
$N^{(S)}_{m}$ is the total data size of clients assigned to server $m$:
$N_{m}^{(S)}=\sum_{i\in\mathcal{U}}n_{i}d_{i,m},$
$D_{m}$ is a distribution parameter:
$D_{m}=\frac{1}{1+|\mathcal{S}|},$
and $\mathbb{I}(TL_{i}<\Theta)$ is an indicator function that equals 1 if the trust level of client $i$ is below the threshold $\Theta$, and 0 otherwise.
By minimizing $B$, we can directly influence and improve the overall performance of the system. Through the detailed derivation above, we have proven Theorem 1. In the proof, we have made standard assumptions such as the strong convexity of the loss function and the Lipschitz continuity of the gradient. We have also taken into account the randomness and error terms during the training process, ultimately deriving an upper bound on the expected difference in the global loss function’s value.